What four consecutive odd integers have a sum of -84
Answer by Alan3354(69443)

(n-2) + n + (n+2) + (n+4) = -84
4n + 4 = -84
4n = -88
n = 22
--> no such numbers
-19 -21 -23 -25 = -88
-17 -19 -21 -23 = -80